A Force of 85.0 N is applied over a distance of 23.4 m to slide a 36.5 kg block across the floor. What is the velocity of the block after the force was applied?
1
Expert's answer
2020-11-10T07:04:21-0500
"W=K\\Fs=0.5mv^2\\85(23.4)=0.5(36.5)V^2\\V=10.4\frac{m}{s}"